# Vision Transformers on the Edge: A Comprehensive Survey of Model Compression and Acceleration Strategies

### v1

@@ 2.2.2 Discussion
A key strength of ViT models lies in their scalability to high parametric complexity; however, this demands significant computational resources and incurs substantial costs. KD offers a way to transfer knowledge into more compact student models, yet challenges remain, particularly in the vision domain. One primary challenge involves the high training costs, as logits-based and hint-based KD methods necessitate extensive GPU memory for the distillation process.
To show the results of the discussed techniques, we divide all the published results into three different Tables (35). The results are divided based on popular CV tasks named image classification (Table 3), object detection (Table 4), and semantic segmentation (Table 5). We observe that most of the papers are tested for image classification tasks, whereas a limited number of papers are evaluated for object detection and semantic segmentation. Table 3 documents all the applied KD in image classification on the ViT backbone. As it is crucial to see the top-1 for classification problems, we summarize these accuracies for teacher and student models from the proposed papers. Moreover, it is essential to calculate the APbox in object detection, and we document the APbox results for both without applying KD and with KD for students model to compare the scenario in Table 4 better. Lastly, Table 5 illustrates the semantic segmentation results for KD methods in ViT. All the current papers in Table 5 used mean intersection over union (mIoU) as a metric to calculate the accuracy. We separately document results with teacher models, student models, and proposed methods to better understand the improvement after applying KD.
@@ 2.3.1 Taxonomy of Quantization Methods
Table 6 gives the overview of different quantization techniques and their pros and cons for computer vision tasks. Table 6 divides the quantization method from the aspects of quantization schemes, quantization approaches, calibration methods, granularity, and other independent techniques. Each aspect has multiple types of quantization techniques applied in different studies. In general, most quantization methods drop accuracy after applying quantization and need finetuning to regain the accuracy.
@@
Calibration Methods Calibration is a needed process of determining the appropriate scaling factors during finetuning methods that map the continuous range of FP values to discrete integer values. Calibration mainly ensures that the range of the quantized values matches the range of the original FP values as closely as possible. There are two types when choosing the range. One is dynamic quantization, and another is static quantization. The weights are quantized statically in dynamic quantization, but activations are quantized dynamically at runtime. Static quantization is quantized post-training. Unlike dynamic quantization, static quantization applies to weights and activations before deploying the model.
Others There are other techniques, such as mixed-precision quantization and hardware-aware quantization. Mixed-precision uses different parts (e.g., channels, layers) of the model that are quantized to different numerical precisions. Unlike uniform quantization, where the entire model is quantized to the same bit-width (like INT8), mixed-precision involves carefully choosing the bit-width for each layer or even each channel within a layer based on their sensitivity and contribution to the final performance of the model. Hardware-aware quantization is another technique that tailors a neural network’s precision reduction process to the specific hardware to deploy on. This technique optimizes the model for the target hardware by adjusting the quantization parameters to match the hardware’s operations [59], such as latency and throughput.
Table 6: Pros and cons for different quantization methods in ViT models.
@@
Techniques Description Pros Cons
@@
Fixed step size between
quantization levels
Simple & Fast
Not adaptive to
@@
Adaptive step size based
on data distribution
@@
More complexity Specific
to data distribution
@@
Post-training
quantization(PTQ)  [64, 65]
Adjust the weights without finetuning
Possible with limited data
& Low computational intensive
Accuracy drop &
Limited adaptation
Quantization-aware
training(QAT)  [66, 67, 68]
Incorporates the quantization process into
the forward and backward propagation steps
of the training algorithm
Better accuracy than PTQ
Complex & High
@@
Apply quantization into weights activation
to low precision before deployment
Computationally efficient Accuracy drop
Dynamic quantization
@@
Quantized the weights statically but
activations are quantized dynamically at
runtime based on data being processed
High accuracy
Slow & Require more
resources at runtime
@@
All weights activations within a layer
are quantized using same scale
Simple & sub-optimal
accuracy
Less flexible
Accuracy drop
@@
Each channels quantized with same
scaling factors
Flexible & High accuracy
Complex & High
computational overhead
Others
@@
Quantized with different bit precision
for each layer
Performance Optimization for
low-precision quantization
Complex & need extended
search processes
@@
Less flexible &
complex
@@ 2.3.2 Quantization Techniques for Vision Transformer
Applying quantization in ViT models is quite new in the neural network sectors. ViT consists of multiple layers, including a self-attention mechanism and a feed-forward neural network. As data passes through these layers, different layers learn to focus on different features of the input data. Applying quantization to ViT can be challenging due to its complexity. Additionally, the loss from quantization can significantly impact the self-attention mechanisms, potentially reducing the model’s overall performance. Quantization methods on ViT can be broadly categorized into two main approaches based on their reliance on training or finetuning: PTQ and QAT.
@@ 2.3.2.1 PTQ Techniques for Vision Transformer
Statistic-Based PTQ methods Most of the current PTQ works on ViT follow statistic-based methods. PTQ4ViT [84], one of the first search-based PTQ methods, addressed two key issues with base-PTQ for ViTs: 1) unbalanced distributions after softmax and asymmetric distributions after GELU. 2) traditional metrics are ineffective for determining quantization parameters. To solve the first problem, the authors proposed twin uniform quantization, which quantizes values into two separate ranges. Additionally, to solve the second problem, they introduced a Hessian-guided metric for improved accuracy instead of mean square error (MSE) and cosine distance.
@@
Building on advancements in PTQ on ViT, Ding et al. [85] proposed APQ-ViT solved two problems of the existing quantization techniques. The authors first proposed a unified bottom-elimination blockwise calibration scheme to solve the inaccurate measurement during quantization value calculation for extremely low-bit representation. This blockwise calibration scheme enables a more precise evaluation of quantization errors by focusing on block-level disturbances that impact the final output. For the second challenge, they observed the "matthew-effect" in the softmax distribution, where smaller values shrink further, and larger values dominate. However, the existing quantizers ignored the mattew-effect of the softmax function, which costs information loss from the larger values. In response, the authors proposed matthew-effect preserving quantization (MPQ) for Softmax to maintain the power-law character to solve the second limitation, ensuring balanced information retention during quantization. Additionally, Liu et al. proposed NoisyQuant [86], where they focused on adding a noisy bias to each layer to modify the input activation distribution before quantization reduced the quantization error. The noisy bias is a single vector sampled from a uniform distribution. The authors removed the impact of noisy bias after the activation-weight multiplication in the linear layer with a denoising bias so that the method could retrieve the correct output. Surprisingly, the experiment showed that adding a noisy bias improved top-1 accuracy compared to the PTQ4ViT [84] on the ViT-B, DEiT-B, and Swin-S model.
@@
Figure 4: Comparison of using full precision Softmax and log-int-softmax in quantized MSA inference in FQ-vit  [14]
@@
However, recent advancements in statistic-based PTQ for ViTs have moved beyond converting FP32 precision (dequantization) during inference, pioneering integer-only fully quantized methods [14, 87, 15]. Moreover, Lin et al. first introduced fully quantized PTQ techniques named FQ-ViT [14] leveraging the power-of-two factor (PTF) method to minimize performance loss and inference complexity. To solve the non-uniform distribution in attention maps and avoid the dequantizing to FP32 before softmax, they proposed log-int-softmax (LIS) replacing softmax. Additionally, they streamline inference further using 4-bit quantization with the bit-shift operator. Figure 4 (left) illustrated the traditional approach (left) where the traditional approach dequantized INT8 query (Q) and key (K) matrices to FP before softmax, re-quantizing afterward for attention computations. In contrast, the proposed method (Figure 4 (right)) introduced matrix multiplication followed by integer-based exponential (i-exp). The authors then utilized $Log_{2}$ quantization scale in the softmax function and converted the MatMul to BitShift between the quantized attention map and values (V). This fully integer workflow, including LIS in INT4 format, significantly reduces memory usage while maintaining precision. Extended from FQ-ViT, Li et al. introduced I-ViT [87], the first integer-only PTQ framework for ViT, enabling inference entirely with integer arithmetic and bit-shifting, eliminating FP operations. In this framework, the authors utilized an integer-only pipeline named dyadic anthemic for non-linear functions such as dense layers. In contrast, non-linear functions, including softmax, GELU, and LayerNorm, were approximated with lightweight integer-based methods. The key contribution of this work is that Shiftmax and ShiftGEU replicated the behavior of their FP counterparts using integer bit-shifting. Despite I-ViT’s reduction in bit-precision for parameters and its emphasis on integer-only inference, it retained its accuracy. For example, when I-ViT applied to DeiT-B, it achieved 81.74% top-1 accuracy with 8-bit integer-only inference, outperforming I-BERT [88] by 0.95% (see Table 7).
@@
However, the current studies consider quantizers and hardware standards always antagonistic, which is partially true. RepQ-ViT [15] decouples the quantization and inference process to explicitly bridged via scale reparameterization between these two steps. The authors applied channel-wise quantization for the post-LayerNorm activations to solve the interchannel variations and $\log\sqrt{2}$ quantization for the post-softmax activations. In the inference, the reparameterized the layer-wise quantization and $\log 2$ quantization with minimal computational cost for respective activations. Using integer-only quantization for all layers lessened the computational cost dramatically and made them highly suitable for edge devices.
@@
Additionally, recent studies have further advanced statistics-based PTQ for ViTs by incorporating mixed-precision techniques. Liu et al.  [89] first explored a mixed precision PTQ scheme for ViT architectures to reduce the memory and computational requirements. The authors estimated optimal low-bit quantization intervals for weights and inputs, used ranking loss to preserve self-attention order, and analyzed layer-wise quantization loss to study mixed precision using the L1-norm [90] of attention maps and outputs. Using calibration datasets from CIFAR-10 [48], ImageNet-1k [51], and COCO2017 [91], their method outperformed percentile-based techniques[73] by 3.35% on CIFAR-10 with ViT-B model. Recently. Tai et al. [92] and Ranjan et al. [93] both extended the mixed precision PTQ techniques on ViT. MPTQ-ViT [92] utilized the smoothQuant [94] with bias term (SQ-b) to address the asymmetry in activations, reducing clamping loss and improving quantization performance. The authors proposed a search-based scaling factor ratio (OPT-m) to determine the quantization parameters. Later, they incorporate SQ-b and OPT-m to propose greedy mixed precision PTQ techniques for ViT by allocating layer-wise bit-width. Additionally, Ranjan et al. [93] proposed LRP-QViT [93], an explainability-based approach by assessing each layer’s contribution to the model’s predictions, guiding the assignment of mixed-precision bit allocations based on layer importance. The authors also clipped the channel-wise quantization to eliminate the outliers from post-LayerNorm activations, mitigating severe inter-channel variations and enhancing quantization robustness. Zhong et al. proposed ERQ [95] to mitigate the error arising during quantization from weight and activation quantization separately. The authors introduced activation quantization error reduction to reduce the activation error, which is like a ridge regression problem. The authors also proposed weight quantization error reduction in an interactive approach by rounding directions of quantized weight coupled with a ridge regression solver.
Learning-Based PTQ Methods While most current PTQ methods for ViTs are statistic-based, there are only a few that utilize learning-based approaches. Existing PTQ methods for ViTs face challenges with inflexible quantization of post-softmax and post-GELU activations, which follow power-law-like distributions. To solve this problem, Wu et al. proposed Adalog [96]. The authors optimized the logarithmic base to better align with the power-law distribution of activations while ensuring hardware-friendly quantization. The authors applied their proposed methods to post-softmax and post-GELU activations through bias reparameterization. Additionally, a fast progressive combining search strategy is proposed to efficiently determine the optimal logarithm base, scaling factors, and zero points for uniform quantizers. Moreover, a recently proposed by Ramachandran et al. named CLAMP-ViT [97] adopted a two-stage approach between data generation and model quantization. The authors introduced a patch-level constrastive learning scheme to generate meaningful data. The authors also leveraged contrastive learning in layer-wise evolutionary search for fixed and mixed-precision quantization to identify optimal quantization parameters. In conclusion, the learned-based PTQ techniques on ViT are largely explored for low-bit quantization.

### v2

@@ 2.2.2 Discussion
A key strength of ViT models lies in their scalability to high parametric complexity; however, this demands significant computational resources and incurs substantial costs. KD offers a way to transfer knowledge into more compact student models, yet challenges remain, particularly in the vision domain.
From the results in Tables (68), we observe a consistent trend: KD techniques can significantly improve the performance of student ViT models. In Table 6, relation-based KD techniques such as Co-advise [57] demonstrate noticeable top-1 accuracy improvements on the ImageNet-1k dataset, emphasizing the effectiveness of capturing inter-sample relationships. Feature-based techniques like MiniViT [59] achieve considerably higher top-1 accuracy than manifold distillation [58] by leveraging intermediate-level information such as attention and hidden states. In contrast, while logit-based methods offer more straightforward supervision through output alignment, their performance improvements tend to be more subtle and, in some cases (e.g., DearKD [56]) demonstrate slight accuracy degradation.
In object detection tasks (see Table 7), KD techniques yield consistent gains in APbox, indicating successful knowledge transfer in detection pipelines. Although the improvements are smaller than in classification tasks, they still affirm the benefit of distillation in dense prediction contexts. For segmentation tasks (see Table 8), the application of KD leads to substantial mean intersection over union (mIoU) improvements. Additionally, we observe that the majority of existing KD approaches for ViT models have been proposed for classification tasks. At the same time, significantly fewer works have explored their applicability to object detection and segmentation tasks. This highlights a promising avenue for future research, particularly as vision-driven domains such as autonomous driving and medical imaging continue to gain momentum and demand high-performance yet efficient models.
@@ 2.3.1 Taxonomy of Quantization Methods
Table 9 categorizes quantization techniques based on their role in the quantization pipeline, including when quantization is applied (timing), how it is performed (scheme), how scaling factors are chosen (calibration), the granularity of quantization, and precision Strategies strategies.
@@
Table 9: Categorization of quantization techniques across five key aspects: schemes, approaches, calibration methods, granularity levels, and deployment strategies.
@@
Techniques Advantages Limitation
@@
Simple implementation &
fast inference
Inflexible to
@@
Complex implementation &
data-dependent
@@
PTQ
 [14, 70]
No retraining required &
low overhead
Accuracy drop
QAT
 [71, 72, 73]
Preserves accuracy &
robust
Requires retraining &
@@
Lightweight Accuracy drop
Dynamic
@@
Adaptive to
input variations
Runtime overhead
@@
Simple implementation
Suboptimal precision
for sensitive layers
@@
Fine-grained control Computational intensive
Precision
@@
Balances accuracy &
efficiency
Not hardware-friendly
@@
Platform-specific
Calibration Methods Calibration is a needed process of determining the appropriate scaling factors during finetuning methods that map the continuous range of FP values to discrete integer values. Calibration mainly ensures that the range of the quantized values matches the range of the original FP values as closely as possible. There are two types when choosing the range. One is dynamic quantization, and the other is static quantization. The weights are quantized statically in dynamic quantization, but activations are quantized dynamically at runtime. Static quantization is quantized post-training. Unlike dynamic quantization, static quantization applies to weights and activations before deploying the model.
Precision Strategies The precision strategies can be divided into three types: fixed precision, mixed-precision, and hardware-aware precision. In fixed-precision quantization, the entire model is quantized using a uniform bit-width (e.g., INT8), simplifying implementation but potentially leading to suboptimal accuracy for layers more sensitive to quantization. To address this limitation, mixed-precision during quantization has been introduced as a more flexible alternative.. Mixed-precision uses different parts (e.g., channels, layers) of the model that are quantized to different numerical precisions. Unlike fixed precision, where the entire model is quantized to the same bit-width (like INT8), mixed-precision involves carefully choosing the bit-width for each layer or even each channel within a layer based on their sensitivity and contribution to the final performance of the model. Despite these advantages, mixed-precision quantization is often not hardware-friendly due to the overhead of managing multiple precision levels within the same model. Hardware-aware quantization has emerged as a solution that tailors a neural network’s precision reduction process according to the hardware requirement. This technique optimizes the model for the target hardware by adjusting the quantization parameters to match the hardware’s operations [85], such as latency and throughput.
Figure 7: An overview of the quantization framework in PTQ4ViT [66], which integrates twin uniform quantization and a Hessian-guided metric for determining scaling factors. Building on this, APQ-ViT [88] addresses the critical quantization error caused by Hessian-based scaling through a block-wise evaluation.
@@ 2.3.2 Quantization Techniques for Vision Transformer
Applying quantization in ViT models is quite new in the neural network sector. ViT consists of multiple layers, including a self-attention mechanism and a feedforward neural network. As data passes through these layers, different layers learn to focus on different features of the input data. Applying quantization to ViT can be challenging due to its complexity. Additionally, the loss from quantization can significantly impact the self-attention mechanisms, potentially reducing the model’s overall performance. Quantization methods on ViT can be broadly categorized into two main approaches based on their reliance on training or finetuning: PTQ and QAT.
@@ 2.3.2.1 PTQ Techniques for Vision Transformer
Statistic-Based PTQ methods Most of the current PTQ works on ViT follow. statistic-based methods. As illustrated in Figure 7, PTQ4ViT [66] addressed two key issues with base-PTQ for ViTs: 1) Unbalanced distributions after softmax and asymmetric distributions after GELU. 2) Traditional metrics are ineffective for determining quantization parameters. To solve the first problem, the authors proposed twin uniform quantization, which quantized values into two separate ranges. Additionally, to solve the second problem, they introduced a Hessian-guided metric for improved accuracy instead of MSE and cosine distance. While effective in general cases, the Hessian-guided metric struggles in ultra low-bit scenarios (e.g., 2-bit or 4-bit), often missing key quantization errors. APQ-ViT, proposed by Ding et al. [88], addresses this limitation through refined error modeling. The authors first proposed a unified bottom-elimination block-wise calibration for extremely low-bit representation. This block-wise calibration scheme enables a more precise evaluation of quantization errors by focusing on block-level disturbances that impact the final output. For the second challenge, they observed the "matthew-effect" in the softmax distribution, where smaller values shrink further, and larger values dominate. Existing quantizers often overlook this issue, leading to information loss from the larger values. To address this, the authors proposed matthew-effect preserving quantization (MPQ) for softmax to retain its power-law characteristics and ensure more balanced information retention during quantization.
@@
Quantization Error with Noisy Bias Liu et al. proposed NoisyQuant [92], where they focused on adding a noisy bias to each layer to modify the input activation distribution to reduce the quantization error. The noisy bias is a single vector sampled from a uniform distribution. The authors removed the impact of noisy bias after the activation-weight multiplication in the linear layer with a denoising bias so that the method could retrieve the correct output. Surprisingly, the experiment showed that adding a noisy bias improved top-1 accuracy compared to the PTQ4ViT [66] on the ViT-B, DeiT-B, and Swin-S model.
@@
Figure 8: Comparison of using full precision Softmax and log-int-softmax in quantized MSA inference in FQ-ViT [14].
@@
Integer-only PTQ Methods Unlike earlier PTQ approaches that rely on FP32 precision (via dequantization) during inference, recent methods have introduced fully integer-only quantization for ViTs [14, 93, 15]. These advancements eliminate floating-point dependencies, reduce computational overhead, and significantly enhance the suitability of ViTs for deployment on resource-constrained edge devices. Lin et al. first introduced fully quantized PTQ techniques named FQ-ViT [14], leveraging the power-of-two factor (PTF) method to minimize performance loss and inference complexity. To solve the non-uniform distribution in attention maps and avoid dequantizing to FP32 before softmax, they proposed that log-int-softmax (LIS) be replaced by softmax. Additionally, the authors further streamlined inference by employing 4-bit quantization combined with bit-shift operations. Figure 8 (left) illustrates the traditional approach where the traditional approach dequantizes INT8 query (Q) and key (K) matrices to FP before softmax, re-quantizing afterward for attention computations. In contrast, the proposed method (Figure 8 (right)) introduced matrix multiplication followed by integer-based exponential (i-exp). The authors then utilized $Log_{2}$ quantization scale in the softmax function and converted the MatMul to BitShift between the quantized attention map and values (V). This integer-only workflow, including LIS in INT4 format, significantly reduces memory usage while maintaining precision. Extended from FQ-ViT, Li et al. introduced I-ViT [93], the first integer-only PTQ framework for ViT, enabling inference entirely with integer arithmetic and bit-shifting, eliminating FP operations. In this framework, the authors utilized an integer-only pipeline named dyadic anthemic for non-linear functions such as dense layers. In contrast, non-linear functions, including softmax, GELU, and LayerNorm, were approximated with lightweight integer-based methods. The key contribution of this work is that Shiftmax and ShiftGEU replicated the behavior of their FP counterparts using integer bit-shifting. Despite I-ViT’s reduction in bit-precision for parameters and its emphasis on integer-only inference, it retained its accuracy. For example, when I-ViT applied to DeiT-B, it achieved 81.74% top-1 accuracy with 8-bit integer-only inference, outperforming I-BERT [94] by 0.95% (see Table 10).
Algorithm 1 Fast Progressive Combining Search (FPCS)
1:  Input: Coefficients $x,y,z_{1},z_{2},k,p$; a pretrained full-precision model; a set of calibration data $\mathcal{D}_{\text{calib}}$; and the $l$-th layer to be quantized $\phi_{l}$.
2:  Output: Quantization hyperparameters $a^{*},b^{*}$
3:  # The initialization step:
4:  Generate the raw input $X_{l}$ and output $O_{l}$ by $\phi_{l}$ based on $\mathcal{D}_{\text{calib}}$, and compute the percentiles $pct_{0},pct_{0.1},pct_{0.9}$ and $pct_{1}$ by [78].
5:  Compute the uniform partition of the first and second hyperparameters as $\mathcal{A}=\{pct_{0.1}+i\cdot\tau_{A}|i=0,\cdots,x\}$ and =$\mathcal{B}=\{pct_{0.9}+j\cdot\tau_{B}|j=0,\cdots,y\}$ with $\tau_{A}=(pct_{0}-pct_{0.1})/x$ and $\tau_{B}=(pct_{1}-pct_{0.9})/y$.
6:  Generate the candidate set $C_{0}$ as the Cartesian product of $\mathcal{A}$ and $\mathcal{B}$: $C_{0}=\mathcal{A}\times\mathcal{B}$.
7:  # The progressive searching step:
8:  for $i=0,\cdots,p$ do
9:     # The coarse searching step:
10:     Construct the subset $C^{\prime}\subset C_{i}$ by selecting the partitions that have the top-$k$ smallest quantization loss.
11:     # The expanding step:
12:     Update the intervals for fine partitions: $\tau_{A}:=\tau_{A}/(2\cdot z_{1})$, $\tau_{B}:=\tau_{B}/(2\cdot z_{2})$.
13:     Update the candidate set with fine partitions: $C_{i+1}=\{(a+i\cdot\tau_{A},b+j\cdot\tau_{B})|(a,b)\in C^{\prime},i=-z_{1},\cdots,z_{1};j=-z_{2},\cdots,z_{2}\}$.
14:  end for
15:  return $(a^{*},b^{*})\in C_{p}$ with the smallest quantization loss.
@@
Decoupling Quantization for Edge-Efficient Inference The current studies consider quantizers and hardware standards to be always antagonistic, which is partially true. RepQ-ViT [15] decouples the quantization and inference process to explicitly bridge via scale reparameterization between these two steps. The authors applied channel-wise quantization for the post-LayerNorm activations to solve the inter-channel variations and $\log_{\sqrt{2}}$ quantization for the post-softmax activations. During inference, the proposed method efficiently reparameterized layer-wise $\log_{2}$-based quantization to reduce the computational overhead associated with activation processing. Using integer-only quantization for all layers lessened the computational cost dramatically and made them highly suitable for edge devices.
@@
Mixed-precision PTQ methods Recent studies have further advanced statistics-based PTQ for ViTs by incorporating mixed-precision techniques. Liu et al. [70] first explored a mixed precision PTQ scheme for ViT architectures to reduce the memory and computational requirements. The authors estimated optimal low-bit quantization intervals for weights and inputs, used ranking loss to preserve self-attention order, and analyzed layer-wise quantization loss to study mixed precision using the L1-norm [95] of attention maps and outputs. Using calibration datasets from CIFAR-10 [52], ImageNet-1k [45], and COCO2017 [96], their method outperformed percentile-based techniques[78] by 3.35% on CIFAR-10 with ViT-B model. Recently. Tai et al. [97] and Ranjan et al. [80] both extended the mixed precision PTQ techniques on ViT. MPTQ-ViT [97] utilized the smoothQuant [98] with bias term (SQ-b) to address the asymmetry in activations, reducing clamping loss and improving quantization performance. The authors proposed a search-based scaling factor ratio (OPT-m) to determine the quantization parameters. Later, they incorporate SQ-b and OPT-m to propose greedy mixed precision PTQ techniques for ViT by allocating layer-wise bit-width. Additionally, Ranjan et al. [80] proposed LRP-QViT [80], an explainability-based approach by assessing each layer’s contribution to the model’s predictions, guiding the assignment of mixed-precision bit allocations based on layer importance. The authors also clipped the channel-wise quantization to eliminate the outliers from post-LayerNorm activations, mitigating severe inter-channel variations and enhancing quantization robustness. Zhong et al. proposed ERQ [99] to mitigate the error arising during quantization from weight and activation quantization separately. The authors introduced activation quantization error reduction to reduce the activation error, which is like a ridge regression problem. The authors also proposed weight quantization error reduction in an interactive approach by rounding directions of quantized weights coupled with a ridge regression solver.
Learning-Based PTQ Methods While most current PTQ methods for ViTs are statistic-based, there are only a few that utilize learning-based approaches. Existing PTQ methods for ViTs face challenges with inflexible quantization of post-softmax and post-GELU activations, which follow power-law-like distributions. One of the major challenges in learning-based approaches is efficiently determining the optimal quantization parameters, such as scaling factors. Adalog [100] introduced a fast progressive combining search (FPCS) strategy, which is designed to search the hyperparameter space effectively while maintaining linear complexity. The FPCS algorithm comprises two stages: initialization and progressive search. The latter performs a coarse search to select top candidate pairs, followed by an expansion step that locally refines the search. The complete procedure is outlined in Algorithm 1. On the other hand, CLAMP-ViT [101] leveraged contrastive learning in layer-wise evolutionary search to identify optimal quantization parameters.
As illustrated in Figure 10, Adalog explored the adaptive $\log_{2}$ and $\log_{\sqrt{2}}$ to better align with the power-law distribution of activations while ensuring hardware-friendly quantization. The adaptive logarithm quantizers are applied to post-softmax and post-GELU activations through bias reparameterization. CLAMP-ViT [101] adopted a two-stage approach between data generation and model quantization. As illustrated in Figure 10, stage 1 is to generate semantically rich and meaningful images. The stage 1 loss is used to adaptively refine the synthetic dataset in a self-supervised manner, forming a feedback loop with the generator. The refined synthetic images generated in Stage 1 are passed to Stage 2. Here, an evolutionary search is conducted on the quantized model to identify optimal layer-wise quantization parameters (e.g., bit-widths, scale factors). This search aims to minimize the Stage 2 loss, which reflects the fitness of quantized representations compared to the FP32 model. However, learning-based PTQ remains largely unexplored for low-bit quantization. Incorporating NAS to discover optimal quantization parameters automatically presents a promising future research direction instead of an evolutionary search.
Figure 9: The Overview of the AdaLog quantization method. AdaLog is applied to post-Softmax and post-GELU activations, with bias parameterization and a search strategy [100].
Figure 10: The Overview of the Clamp-ViT [101]. Stage 1 corresponds to steps ①–②, Stage 2 corresponds to steps ③–④, and the iterative refinement process is denoted by step ⑤.